OpenStudy (rock_mit182):

For the following right triangle, find the side length . Round your answer to the nearest hundredth.

OpenStudy (rock_mit182):



OpenStudy (rock_mit182):

i need the exact value to be verified: c = 11,168 ?

OpenStudy (rock_mit182):

please help quickly

OpenStudy (rock_mit182):

am i Right? is all i need to know the time is running

OpenStudy (wolf1728):

x² = 16² - 11² x² = 256 - 121 x² = sq root (135) x = 11.6189500386

OpenStudy (wolf1728):

Use a period and NOT a comma

OpenStudy (rock_mit182):

why is that

OpenStudy (rock_mit182):

so is 11.168 rounded ?

OpenStudy (wolf1728):

Yees

OpenStudy (wolf1728):

Actually 11.619 rounded

OpenStudy (rock_mit182):

oh men i can't fail im going to trust in you

OpenStudy (wolf1728):

11.619 DEFINITELY

OpenStudy (rock_mit182):

thanks a lot I'm going to answer

OpenStudy (wolf1728):

Go ahead The reason I was picky about the period and the comma was that if your test was being automatically scored - the comma would have made your first answer wrong

OpenStudy (rock_mit182):

oh i see

OpenStudy (wolf1728):

okay By the way here is a Pythagoren Theorem calculator http://www.1728.org/pythgorn.htm enter those numbers and see what you get

OpenStudy (wolf1728):

You'll get 11.619 :-)